OpenStudy (anonymous):

PLEASE HELP!! :D Solve by graphing. Show the work involved regardless of the method you use to graph, such as your table of values and the work involved in plugging numbers into the equations, or the slope and y-intercept if you use that method to graph. y=x+3 y= 3/2x-2

OpenStudy (anonymous):

First off, slope (m) is equal to rise/run. Rise means how far you go up or down and run is how far you go right or left. y=x+3 has an y intercept of three and a slope of 1. (y=mx+b, b is the y intercept and m is the slope. You start at 3 and go up 1 right 1 each time. y= 3/2x-2 has a slope of 3/2 and a y intercept of -2. You plot these points starting at -2, going up 3 and over 2 each time. Do you need the graph and table of values as well or can you do it?

OpenStudy (anonymous):

I think I can Do it Thank you so much!! :D

OpenStudy (anonymous):

No problem. Let me know if there's anything else. :3

OpenStudy (anonymous):

Actually What would the table values be? (x,y)

OpenStudy (anonymous):

For the first equation I will go from -7 to 7. -4=-7+3 (-7,-4) -3=-6+3 (-6,-3) -2=-5+3 (-5, -2) -1=-4+3 (-4,-1 0=-3+3 (-3,0) 1=-2+3 (-2,1) 2=-1+3 (-1,2) 3=0+3 (0,3) 4=1+3 (1,4) 5=2+3 (2,5) 6=3+3 (3,6) 7=4+3 (4,7) 8=5+3 (5,8) 9=6+3 (6,9) 10=7+3 (7,10) ____________________________________________________________________________________ Instead I just googled table calculator and got this.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem.